OpenStudy (anonymous):

Medal Evaluate by using the Order of Operations 24 / 3^2/8-5 - (-5)

jagr2713 (jagr2713):

can you put it in the equation box please

OpenStudy (anonymous):

i will try lol

OpenStudy (anonymous):

\[24 \div \frac{ 3^2 }{ 8-5 } - (-5)\]

jagr2713 (jagr2713):

Thanks

OpenStudy (anonymous):

your welcome

jagr2713 (jagr2713):

Opp sorry gtg do something for 10mins let me ask someone who can help @robtobey @Kainui @sleepyjess

OpenStudy (anonymous):

ok thank you

OpenStudy (sleepyjess):

hello! :) First off, do you know the order of operations?

OpenStudy (anonymous):

bedmas

OpenStudy (sleepyjess):

PEMDAS

OpenStudy (anonymous):

no bedmas

OpenStudy (sleepyjess):

what does each letter stand for?

OpenStudy (anonymous):

brackets exponents division multiply addition subtraction

OpenStudy (sleepyjess):

hmm... interesting

OpenStudy (sleepyjess):

are you in the US?

OpenStudy (anonymous):

no i am in canada

OpenStudy (sleepyjess):

We use PEMDAS here Parentheses Exponents Multiply Divide Addition Subtraction

OpenStudy (sleepyjess):

But anyways, now that we got that cleared up, what is in brackets?

OpenStudy (anonymous):

Refer to the attachment from Mathematica v9.

OpenStudy (anonymous):

(-5)

OpenStudy (sleepyjess):

yep, so - (-5) is the same as +5, so now we get \(24 \div \dfrac{ 3^2 }{ 8-5 } +5\)

OpenStudy (anonymous):

yes now exponents 3^2 = 9

OpenStudy (sleepyjess):

Yep, so now \(\sf 24 \div \dfrac 9{8-5} + 5\) Then 8-5 on the bottom would leave 3 \(\sf 24 \div \dfrac 9{3} + 5\) What do you think now?

OpenStudy (anonymous):

24/ 1 divide by 9/3 + 5

OpenStudy (sleepyjess):

What is 9/3?

OpenStudy (anonymous):

3

OpenStudy (sleepyjess):

Yep, so 24/3 now

OpenStudy (anonymous):

8

OpenStudy (sleepyjess):

let's hope this goes through 8+5 = final answer

OpenStudy (anonymous):

13

OpenStudy (sleepyjess):

yep :)

OpenStudy (anonymous):

so the answer is 13